SpaceRef

SpaceRef


SOHO Daily Meeting Minutes for Thursday 26 June 2003

Status Report From: Goddard Space Flight Center
Posted: Thursday, June 26, 2003

ANNOUNCEMENTS

* We expect to lose the high gain antenna link late this week for ~20 days. We have daily passes on 34 meter stations for monitoring and commanding of the spacecraft. The 34 meter schedule for this week is included below. Since the times are very limited, we'd like to request that ALL TEAMS BE AVAILABLE (at EOF or by phone) during the passes, in case your instrument comes up in an unexpected state. When downlink only 34 meter stations are scheduled, a 26 meter station will be used for uplink (i.e. all 34 meter passes will have commanding capability).

Week 26 34-Meter Passes

                   BOT   EOT
                 --------------
   Thurs   DSS-54 15:30-17:45 UT
   Fri     DSS-54 13:00-17:45 UT
   Sat     DSS-24 17:35-20:35 UT
   Sun     DSS-54 13:00-17:30 UT

* Current RHESSI default target is AR10391. When selecting this target to maximize coordination w/others, use campaign number 6850.

* SPWG on Friday in the EOF at 10am.

* CNE network outage from 6-8pm on Wednesday 2 July. Internet access, e-mail services and building to building traffic will be intermittent.

* 34m stations for next week:

  Monday:    D66 05:20-16:55 26-meter
             D63 12:00-16:00

Tuesday: D54 10:15-14:15

Wednesday: D66 05:40-17:00 26-meter D65 13:00-17:00 D/L

Thursday: D34 00:20-04:20 D24 15:00-19:00

Friday: D24 15:45-19:45

Saturday: D24 17:05-21:05

Sunday: D34 01:50-04:45 D66 06:00-18:30 26-meter

FOT REPORT

Spacecraft Status: Nominal: HGA Position Monitoring Disabled, HGA Nominal Control Disabled

Spacecraft Anomalies: None

Accomplished Activities: NONCOHO_03176

Planned Activities: HGA Test Activities

Upcoming Operations: None

Ground Anomalies:

176/1322 9 minute TM/CMD outage. D16. TCP Failure. TM is recoveable from SSR.

176/1430 61 minutes degraded data. D16. RFI. 1516 - 1531 is recoverable.

SOLAR STATUS

Web Page for Planning: http://www.bbso.njit.edu/arm

Magnetic Maps and Heliospheric Forecasts: http://www.lmsal.com/forecast/

INSTRUMENT STATUS

CDS: Safed - doors closed.

UVCS: Safed - door closed.

LASCO: Safed - doors closed.

EIT: Bakeout.

MDI: Nominal. Will safe when we get to low rate.

TRACE: Nominal. Observing AR10391 & AR10387, full disk mosaics in 171, 284, 1600 & WL.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter